Find the maximum value and minimum value of f(x,y,z)=xyz subject to x+y+z=1 x≥0; y≥0; z≥0


1
Expert's answer
2022-03-04T09:36:12-0500

We are asked for the extreme values of "f" subject to the constraint "g(x,y,z)=x+y+z=1."

Using Lagrange multipliers, we solve the equations "\nabla f(x,y)=\lambda\nabla g(x, y)" and "g(x,y)=1,"

which can be written as


"f_x=\lambda g_x, f_y=\lambda g_y,f_z=\lambda g_z, x+y+z=1"

"yz=\lambda, xz=\lambda,xy=\lambda, x+y+z=1"


"yz=xz=xy"

"yz=xz=>z=0\ or\ x=y"

If "z=0"


"xy=0, x+y=1"

We have "x=0,y=1, z=0" or "x=1,y=0, z=0."


If "z\not=0," then "x=y."

If "x=y=0," then "z=1."

If "x=y\not=0," then "z=x=y=1\/3."


"f(0,1,0)=f(1,0,0)=f(0,0,1)=0"

"f(\dfrac{1}{3},\dfrac{1}{3},\dfrac{1}{3})=\dfrac{1}{3}(\dfrac{1}{3})(\dfrac{1}{3})=\dfrac{1}{27}"

Therefore the maximum value of "f" on the plane "x+y+z=1" is

"f(\dfrac{1}{3},\dfrac{1}{3},\dfrac{1}{3})=\dfrac{1}{27},"

and the minimum value is

"f(0,1,0)=f(1,0,0)=f(0,0,1)=0"